OpenStudy (anonymous):

I am solving a quadratic equation and I and my a term is 22r^2, b term is 19r and c is 4. How would I factor this?

OpenStudy (anonymous):

is the equation\[22r ^{2} + 19r + 4 = 0 ???\]

OpenStudy (anonymous):

Yes, but it doesn't say what it equals.

OpenStudy (anonymous):

ok so u want to factorise it.. proceed as follows here a = 22 and c=4 so ac = 88 we have to split 88 into two factors A and B so tha AB = 88 and A+B = 19 88 = 2x44 , 4x22, 8x11 out of these combinations, 8x11 fulfills our requirement 8x11 = 88 and 8+11 = 19 so 22r^2 + 19r + 4 = 22r^2 + 8r + 11r + 4 = 2r(11r +4) +1(11r +4) = (2r+1)(11r+4) this way you get two factors of the give quadratic equation

OpenStudy (anonymous):

If answer is clear, pls click GOOD ANSWER button for me.... :))

OpenStudy (anonymous):

????????????